Part2 Option No. 2-Investor Capital Semih has found an investor who is willing to provide him with 500,000 TL in exchange for 30% ownership, which means he will take 30% of the annual profits. The investor generally seeks a 12% ROI for his investments. Lets say Semih anticipates that he will have future profits of 200,000 TL peryear for the forseeable future. What is the Present Value of the investment for the investor? Based on the investment amount, what is the total valuation of Semihs company? (Hint: if 30%-500,000 TL, then 100%-?) b) Profit Projections Year 100,000.00 25,000.00 156,000.00 195,000.00 244,000.00 305,000.00 81,000.00 477,000.00 596,000.00 745,000.00 10
Answer #1

3a) Profit for investor=30%*200000=60000
Present value=profit/roi
=60000/12%=500000
b)Total valuation=500000*(100%/30%)=1666667